# RED-PSM: Regularization by Denoising of Partially Separable Models for Dynamic Imaging

### v3

@@ Abstract
Dynamic imaging addresses the recovery of a time-varying 2D or 3D object at each time instant using its undersampled measurements. In particular, in the case of dynamic tomography, only a single projection at a single view angle may be available at a time, making the problem severely ill-posed. In this work, we propose an approach, RED-PSM, which combines for the first time two powerful techniques to address this challenging imaging problem. The first, are partially separable models, which have been used to efficiently introduce a low-rank prior for the spatio-temporal object. The second is the recent Regularization by Denoising (RED), which provides a flexible framework to exploit the impressive performance of state-of-the-art image denoising algorithms, for various inverse problems. We propose a partially separable objective with RED and a computationally efficient and scalable optimization scheme with variable splitting and ADMM. Theoretical analysis proves the convergence of our objective to a value corresponding to a stationary point satisfying the first-order optimality conditions. Convergence is accelerated by a particular projection-domain-based initialization. We demonstrate the performance and computational improvements of our proposed RED-PSM with a learned image denoiser by comparing it to a recent deep-prior-based method known as TD-DIP. Although the main focus is on dynamic tomography, we also show performance advantages of RED-PSM in a cardiac dynamic MRI setting.
11footnotetext: This research was supported in part by Los Alamos National Labs under Subcontract No. 599416/CW13995.
@@ I-A Proposed Approach
We propose an approach, RED-PSM, which combines for the first time two powerful techniques to address this challenging imaging problem. The first, are partially separable models [9, 10, 11, 12, 13, 14, 15, 16], which have been used before to efficiently introduce a low-rank prior for the spatio-temporal object. The second is the recent Regularization by Denoising (RED) [17], which provides a flexible framework to exploit the impressive performance of state-of-the-art image denoising algorithms, for various inverse problems.
We propose a partially separable objective with RED and a computationally efficient and scalable optimization scheme with variable splitting and ADMM.
@@
We demonstrate the performance and computational improvements of our proposed RED-PSM with a learned image denoiser by comparing it to a recent deep-prior-based method known as TD-DIP [18], and to spatial and spatiotemporal total variation (TV) regularized versions of PSM.
Although the main focus is on dynamic tomography, we also show performance advantages of RED-PSM in a cardiac dynamic MRI setting.
@@
Our motivation for using the original RED formulation in RED-PSM rather than more recent formulations and analyses (e.g., as in [22] and the references therein) is based on its simple gradient expression, and explicit regularizer expression facilitating implementation and theoretical analysis. The objective is to demonstrate the large improvement in performance that this simple learned regularizer can bring to the PSM scheme, and the improvement over much more complex and slower state-of-the-art methods. Recent variations on RED such as in [22], and PnP methods (as covered in [20]) have improved over the original methods in reconstruction quality in various static imaging problems and/or in theoretical guarantees. However, theoretical guarantees for the case of a non-convex data fidelity term have been provided recently for only one such non-convex scenario - phase-retrieval [23]. As this is the first work to extend RED to spatio-temporal imaging with a PSM model, and to provide a convergence guarantee for the method in the non-convex bilinear scenario, we chose to limit our attention in this paper to the original RED version [17, 24].
I-B Previous Work on Dynamic Imaging
Previous work [25, 26] on time-sequential sampling of bandlimited spatio-temporal signals provided an optimal view angle order for the scan and theoretical guarantees for unique and stable reconstruction. However, the approach is limited by its bandlimitedness assumptions.
Several methods [27, 28] recover the underlying motion field from projections using the Radon transform properties. However, these methods require the object information a priori. Other algorithms [29, 30, 31] alternate between estimating the motion field and the time-varying object. However these methods assume the total density to be preserved over time, which may be a limiting assumption. For instance, imaging a fixed slice of a 3D time-varying object under compression may violate this assumption due to cross-slice motion.
On the other hand, a model-based iterative reconstruction method [32] for dynamic synchrotron CT incorporates a spatiotemporal Markov random field object prior, and models measurement imperfections, and noise characteristics. An interlaced projection acquisition scheme is also proposed.
Another method [33] proposes a multi-agent consensus equilibrium [34]-based technique using multiple deep denoisers for different spatial directions for dynamic cone beam CT.
In dynamic tomography and MRI, partially separable models (PSM) with a particular low-rank form partially separating the spatial and temporal structure have been used to represent the underlying object [11, 12, 13, 14, 15, 16] and the time-sequential projections [9, 10]. The projection-domain PSM in [9, 10] carries the PSM to the projection domain using the harmonic representation of projections, and provides uniqueness and stability analysis for the problem. In spite of this advantage of the projection domain PSM over its object-domain counterparts, the performance of this approach is still limited by the null space of the measurement operator.
A different PSM-based method for dynamic MRI [35] has a hybrid objective with a PSM model in the data fidelity term and uses the object in a simple $l_{1}$-norm sparsity penalty, with an additional term to penalize the mismatch between the PSM and the object. The method uses a fixed temporal basis initialization with the assumption of achieving temporal Nyquist sampling rate which is not possible for the time-sequential scenario considered in this work. It does not use an initial guess for the spatial basis functions.
Other PSM-based methods for dynamic MRI [35] and tomography [36] impose the low-rank structure as a soft-constraint with hybrid objectives. The dMRI method [35] uses PSM in the data fidelity term, and uses the object in a basic $l_{1}$-norm sparsity penalty, with an additional term to penalize the mismatch between the PSM and the object. The method does not use a guided initial guess for the object or for the PSM basis functions. The dynamic tomography technique [36] instead uses the object in both the data fidelity term and in a simple total-variation regularizer, with a penalty for object-PSM mismatch, thus providing additional flexibility in the variational model. The method employs a fast initialization scheme using the projection domain PSM algorithm [9, 10].
A recent dMRI method [37] partitions the dynamic object into temporal subsets each lying on a low-dimensional manifold, approximated by its tangent subspace. As a result, the object is represented by a temporally partitioned PSM. A simple sparsity penalty is used for the spatial representation. The spatial basis functions and linear transformations aligning the temporal subsets are determined by optimizing a non-convex objective using an ADMM algorithm, without a convergence analysis. The subset temporal bases are estimated separately from training data, which, in some applications, may not be available. The temporal partitioning uses a heuristic criterion applied to an initial reconstruction.
@@ I-B Previous Work on Dynamic Imaging
In an earlier dynamic tomography method [38], the object is decomposed into a low-rank and a sparse component, and the low rankness is enforced implicitly using the nuclear norm. The method assumes the object as static for a set of projections. Other works perform various similar decompositions for dMRI [39, 40, 41, 42]. A related fast dMRI method [43] decomposes the object representation into three components: the mean signal, a low-rank PSM, and the residual, which is assumed to be sparse in the Fourier domain. Then, each component is estimated in order. Although benchmark competing methods are either faster or more accurate on some of the diverse data sets used for the comparison in [43], the method provides the lowest reconstruction error and runtimes when these quantities are averaged over these data sets.
@@
Recently, object-domain deep image prior (DIP)-based algorithms [18, 44] have been proposed for dynamic MRI. Providing impressive results, DIP-based algorithms such as [18] suffer from overfitting and usually require handcrafted early stopping criteria during the optimization of generator network parameters. To overcome these drawbacks, [44] includes regularization constraining the geodesic distances between generated objects in proportion to the distances between their latent representation. However, this requires the computation of the Jacobian of the generator network at each iteration of the update of the weights, and significantly increases the computational load and run time.
@@
A different approach [45] combines partially separable and generative models. It employs spatial and temporal regularization by penalizing the generator weights and a particular initialization scheme for the inputs to the generator networks. Although it combines the PSM with the recent DIP framework, this method has the following limitations, which are overcome by the proposed approach: (i) The spatial generator is an artifact removal network taking the full-sized spatial basis functions as input. Since this prevents a patch-based implementation, it may be computationally problematic for high-resolution 3D+temporal settings. (ii) The CNN prior for natural images may not be useful as a prior for the individual spatial basis functions, since the least-squares optimal spatial basis functions are the left singular vectors of the complete object, and as such may not have the structure of natural images. (iii) As with the other DIP-based methods, if the additional penalty on the generator parameters is insufficient, this method can be prone to overfitting.
@@
Recently, a neural field-based technique [46] was proposed for dynamic photoacoustic tomography (PACT). The authors use a neural field to represent the dynamic object with a considerably reduced number of parameters. Then, they combine this representation with a simple total variation (TV)-based regularizer and compare the method with the full-rank representation of the object and TV or low-rank inducing priors. Despite providing an efficient representation for the dynamic object, the method overperforms its comparison benchmarks only in some of the view-per-frame settings.
@@ I-C Contributions
  1. 1.
    To the best of our knowledge, RED-PSM is the first PSM-based approach to dynamic imaging, that uses a pre-trained, learned (RED-based [17]) spatial prior.
  2. 2.
    We are not aware of any prior work that uses RED with an explicit low-rankness constraint. In RED-PSM we are able to achieve a parsimonious representation for the dynamic object using PSM while incorporating the RED prior, reducing the data requirements further.
  3. 3.
    Unlike supervised methods for spatio-temporal imaging [47, 48, 49, 50, 51] that learn a spatio-temporal model from training data, RED-PSM does not require ground-truth spatio-temporal training data, which is often not available or expensive to acquire. Thus RED-PSM offers the best of both worlds: a learned spatial model using readily available static training data, and unsupervised single-instance adaptation to the spatio-temporal measurement.
  4. 4.
    A novel and effective ADMM algorithm for the resulting new minimization problem enforces PSM as a hard constraint and incorporates the learned RED spatial regularizer.
  5. 5.
    The method is supported by theoretical analysis, with a convergence guarantee which is of practical importance. First, it provides a mathematical justification to terminate the algorithm after sufficiently many iterations. Second, it eliminates concerns about the solution degrading after too many iterations as in the case of DIP-based methods. While RED has been applied to problems with nonlinear measurement operators, the convergence of RED has been analyzed in this setting only for the phase retrieval problem [23]. The analysis of RED-PSM is the first convergence result of RED in the bilinear, non-convex scenario. Moreover, no convergence analysis is provided for previous PSM works.
  6. 6.
    Compared to a recent DIP-based [18] algorithm, RED-PSM achieves better reconstruction accuracy with orders of magnitude faster run times.
  7. 7.
    To improve and speed up the empirical convergence of RED-PSM, we use a particular fast projection-domain PSM initialization scheme [9, 10] for the spatial and temporal basis functions. The accelerated and reliable convergence with this initialization scheme is important for applications (e.g. bio-medical) that require fast turn-around between acquisition and reconstruction.
  8. 8.
    A version of the approach with a patch-based regularizer is shown to provide almost equivalent reconstruction accuracy. This makes the proposed method conveniently scalable to high-resolution 3D or 4D settings.

### v4

# RED-PSM: Regularization by Denoising of Factorized Low Rank Models for Dynamic Imaging

@@ Abstract
Dynamic imaging addresses the recovery of a time-varying 2D or 3D object at each time instant using its undersampled measurements. In particular, in the case of dynamic tomography, only a single projection at a single view angle may be available at a time, making the problem severely ill-posed. We propose an approach, RED-PSM, which combines for the first time two powerful techniques to address this challenging imaging problem. The first, are non-parametric factorized low rank models, also known as partially separable models (PSMs), which have been used to efficiently introduce a low-rank prior for the spatio-temporal object. The second is the recent Regularization by Denoising (RED), which provides a flexible framework to exploit the impressive performance of state-of-the-art image denoising algorithms, for various inverse problems. We propose a partially separable objective with RED and a computationally efficient and scalable optimization scheme with variable splitting and ADMM. Theoretical analysis proves the convergence of our objective to a value corresponding to a stationary point satisfying the first-order optimality conditions. Convergence is accelerated by a particular projection-domain-based initialization. We demonstrate the performance and computational improvements of our proposed RED-PSM with a learned image denoiser by comparing it to a recent deep-prior-based method known as TD-DIP. Although the main focus is on dynamic tomography, we also show performance advantages of RED-PSM in a cardiac dynamic MRI setting.
Index Terms:
Dynamic imaging, Regularization by denoising, Low rank modeling.
@@ I-A Proposed Approach
We propose an approach, RED-PSM, which combines for the first time two powerful techniques to address this challenging imaging problem. The first technique, are non-parametric factorized low-rank object models, popularized under the name partially separable models (PSMs) [9]. PSM was combined with low-rank matrix recovery in [10, 11, 12], and used since then to represent or motivate a low-rank object prior in tens of works on dynamic imaging. The second technique is the recent Regularization by Denoising (RED) [13], which provides a flexible framework to exploit the impressive performance of state-of-the-art image denoising algorithms, for various inverse problems.
We propose a non-convex partially separable objective with RED for learning-regularized low-rank spatio-temporal object recovery and a computationally efficient and scalable optimization scheme with bi-convex ADMM.
@@
We demonstrate the performance and computational improvements of our proposed RED-PSM with a learned image denoiser by comparing it to a recent deep-prior-based method known as TD-DIP [14], and to spatial and spatiotemporal total variation (TV) regularized versions of PSM with a bi-convex objective.
Although the main focus is on dynamic tomography, we also show the performance advantages of RED-PSM in a cardiac dynamic MRI setting.
@@
Our motivation for using the original RED formulation in RED-PSM rather than more recent formulations and analyses (e.g., as in [18] and the references therein) are its simple gradient expression, and explicit regularizer expression facilitating implementation and theoretical analysis. The objective is to demonstrate the large improvement in performance that this simple learned regularizer can bring to the PSM scheme, and the improvement over much more complex and slower state-of-the-art methods. Recent variations on RED such as in [18], and PnP methods (as covered in [16]) have improved over the original methods in reconstruction quality in various static imaging problems and/or in theoretical guarantees. However, theoretical guarantees for the case of a non-convex data fidelity term have been provided recently for only one such non-convex scenario - phase-retrieval [19]. As this is the first work to extend RED to spatio-temporal imaging with a PSM, and to provide a convergence guarantee for the method in the non-convex bilinear scenario, we chose to limit our attention in this paper to the original RED version [13, 20].
I-B Previous Work
We limit our discussion to dynamic CT and dMRI, although similar or analogous methods have been or can be applied in other dynamic imaging modalities. The same can be said of the method proposed in this paper.
I-B1 Dynamic MRI (dMRI)
In spite of major advances in hardware and signal acquisition methods, MRI has remained a relatively slow modality, posing significant challenges for dynamic imaging. This, coupled with the high flexibility in the acquisition strategy, has motivated extensive work on dMRI, with hundreds of papers on the subject. We highlight some key developments in the algorithm-based dMRI techniques.
Fundamentally, algorithmic approaches to dMRI are based on identifying and exploiting redundancy in the dynamic object, which enable its recovery from an incomplete, and usually sparse set of samples acquired in the joint k-space and time domain – the $k$-$t$-space. Many of the proposed methods have extensions to parallel imaging, which introduces additional such redundancy by additional hardware.
A large set of approaches rely on modeling the support of the object in the domain dual to the $k$-$t$-domain – the ($x$-$f$) domain (the space – temporal frequency domain). These approaches may be classified into three groups. The first involves generic modeling of the ($x$-$f$) support as in reduced field-of-view (FOV) techniques [21, 22, 23]. The second involves sparse and unknown support modeling in a transform domain, and sampling and recovery using the methods of compressed sensing [24, 25, 26, 27]. The third involves adaptive sparse support modeling [28, 29, 30, 31], where the support information is adapted to the dynamics of the imaged slice, and the theory of time-sequential sampling [32, 33] is used to design an optimal $k$-$t$ sampling and reconstruction scheme. Other related approaches model prior information such as motion periodicity and related harmonic structure [34, 35], or the object’s spatial and temporal correlation functions [36].
Since its introduction, the partially-separable model (PSM) [9], with a non-parametric factorized low-rank form separating the spatial and temporal structure, has been used extensively in many works on dMRI to represent the underlying spatio-temporal object. A group of methods recover the PSM in two steps: (a) estimate the temporal subspace using data (e.g., a navigator signal) sampled at low spatial, but high temporal resolution; and (b) estimate the spatial subspace using the entire acquired data set. Such methods include e.g., [9, 37, 38, 39, 40], with followup versions promoting sparsity of the object in a transform domain (e.g., [41]). Other PSM approaches use a low-rank matrix recovery formulation to estimate spatial and temporal components jointly (free of conditions or additional data acquisition enabling a separate reconstruction of the temporal subspace), e.g., [10, 11]), also with versions promoting sparsity [12, 42].
[…]
Another set of methods, including [44, 45, 46, 47], promotes low-rankness of the entire spatio-temporal object matrix or of patches thereof [48, 49] implicitly by using the nuclear norm or Schatten-p functional with $p<1$, often promoting sparsity in a transform domain as well. Methods combining motion estimation and compensation with other priors are also used widely for dMRI, e.g., [50, 51, 52, 53].
@@ I-B1 Dynamic MRI (dMRI)
A different approach in dMRI decomposes the object into the sum of low-rank and sparse components, with the low-rankness encouraged implicitly using nuclear or a Schatten-p functional [54, 55, 56, 57]. A recent related fast method [58] decomposes the object representation into the sum of three components: the mean signal; a low-rank PSM; and a residual sparse in the Fourier domain. Although benchmark competing methods are either faster or more accurate on some of the diverse data sets used for the comparison in [58], the method provides the lowest reconstruction error and runtimes when these quantities are averaged over these data sets.
Recent approaches to dMRI reconstruction have used deep neural networks (DNNs). A majority of these methods, such as [59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], are supervised, requiring a large set of high-resolution artifact-free images for network training. However, acquiring such training data can be difficult for some dMRI applications, and may be impossible for many other dynamic imaging applications, where clinicial dMRI strategies for freezing motion such as ECG synchronization and breath-hold by a cooperative patient, are not available.
@@
The need for a training data set is overcome by recent object-domain deep image prior (DIP)-based DNN algorithms for dMRI [14, 73], which are unsupervised. Providing impressive results, DIP-based algorithms such as [14] suffer from overfitting and usually require handcrafted early stopping criteria during the optimization of generator network parameters. To overcome these drawbacks, [73] includes regularization constraining the geodesic distances between generated objects. However, this requires the computation of the Jacobian of the generator network at each iteration of the update of the weights, significantly increasing the computation and run time.
@@
A different unsupervised approach [74] combines factorized low rank (i.e., partially separable) with generative models. Although it combines the PSM with the recent DIP framework, this method has the following limitations: (i) The spatial generator is an artifact removal network taking the full-sized spatial basis functions as input. Since this prevents a patch-based implementation, it may be computationally problematic for high-resolution 3D+temporal settings; (ii) The CNN structural prior for natural images that is used in the spatial generator may not be useful as a prior for the individual spatial basis functions, since the least-squares optimal spatial basis functions are the left singular vectors of the complete object, and as such may not have the structure of natural images; (iii) As with the other DIP-based methods, if the additional penalties on the generator parameters are insufficient, this method can be prone to overfitting. These various limitations are overcome by the proposed approach.
I-B2 Dynamic Tomography (dCT)
CT is faster than MRI, thanks to acquiring an entire projection at one view angle at once. However, because of the angular scanning required in almost all systems, in spite of significant advances in hardware, dCT is still a relatively slow modality. It is also the problem that motivated much of our work reported in this paper.
In perhaps the earliest algorithmic approaches to dCT, the time-sequential sampling problem in d-CT was formulated and studied in [75], and [76, 77] provided a solution using time-sequential sampling of bandlimited spatio-temporal signals including an optimal view angle order for the scan and theoretical guarantees for unique and stable reconstruction. However, the approach is limited by its bandlimitedness assumptions.
A class of algorithms for dCT are based on modeling the time variation by a motion field, and motion-compensated reconstruction. Many of these algorithms, e.g., [78, 79, 80, 81] alternate between estimating the motion field and the time-varying object. Some algorithms, e.g., [82] use the optical flow (mass continuity) PDE. A different method [83] estimates the motion field separately by evolving a linear elastic deformation by the Navier-Cauchy PDE, assuming known boundary evolution of the object and initial density distribution. However, these methods assume the total mass or density to be preserved over time, which may be limiting assumptions. For instance, flow of a contrast agent into the imaged volume, or imaging a fixed slice of a 3D time-varying object under compression with cross-slice motion may violate either of these assumptions. As a hardware-based, problem-specific approach, optical tracking of fiducial marks has been used for motion correction [84].
Various methods approach dCT reconstruction as Bayesian estimation. These include methods [85, 86, 87, 88] that employ a state-space formulation and use Kalman filter techniques to approximate an MMSE estimate. A different Bayesian method [89] incorporates a spatiotemporal Markov random field object prior, and models measurement imperfections. An iterative algorithm is used to compute a MAP estimate. An interlaced projection acquisition scheme is also proposed.
Several dCT methods motivated by compressed sensing use a total-variation penalty along both spatial and temporal coordinates as in [90], while other methods, e.g., [91] propose efficient sparse representations for the dynamic object. Incorporating low-rank modeling [92] decomposes the object into low-rank+sparse components, promoting the low rankness implicitly using the nuclear norm. The method relies on the object being approximately static for groups of projections, enabling an approximate reconstruction from each such group.
The PSM model has been introduced into dynamic tomography [93, 94] by carrying the PSM to the projection domain using the harmonic representation of projections, and estimating the spatial and temporal basis functions jointly. The work provides a theoretical analysis of uniqueness and stability and the choice of time-sequential angular scan scheme for the problem. Despite the advantages of this approach, its performance is still limited by the null space of the measurement operator.
To help address the underdetermined problem, the method of [95] improves over [93, 94] by combining a PSM with basic spatial regularization by total variation. Like [93, 94] this method too estimates the temporal and spatial PSM components jointly. This makes it applicable to the time-sequential acquisition scenario considered in this paper, where only a single projection is available per time frame. It imposes low rankness as a soft constraint, using a hybrid biconvex objective with a data fidelity term and a TV regularizer expressed in terms of the unfactorized object, and a penalty for object-PSM mismatch. It uses [93, 94] for fast initialization. The method [95] can be considered a precursor to the method of this paper, which combines a hard rank constrained PSM with a sophisticated learned spatial regularizer, and provides theoretical convergence analysis.
Recent methods [96, 97, 98, 99] introduce deep learning into dCT. These methods require an initial reconstruction from the subsets of data from which the motion can be estimated accurately. However, without a periodicity assumption, this is not applicable to the scenarios in this paper, where only a single projection is available at a time. Another method [100] for cone-beam dCT proposes a multi-agent consensus equilibrium [101]-based technique using separate deep denoisers for axial, sagittal, and coronal slices of the object.
@@ I-B2 Dynamic Tomography (dCT)
A deep-learning-based method [102] for dynamic photoacoustic tomography (PACT) introduced a neural field (NF) to represent the dynamic object with fewer parameters, combining it with a simple total variation (TV)-based regularizer. Despite providing an efficient representation for the dynamic object, the method outperforms its comparison benchmarks only in some of the projection-per-frame settings.
@@ I-C Contributions
1. To the best of our knowledge, RED-PSM is the first PSM-based approach to dynamic imaging that incorporates a particular (RED-based [13]) spatial prior that is learned and pre-trained.
2. We are not aware of any prior work that uses RED with an explicit low-rankness constraint. In RED-PSM we achieve parsimonious representation for the dynamic object using PSM, and reduce the data requirements further by incorporating the RED prior.
3. Unlike supervised learning-based methods for spatio-temporal imaging [59, 60, 61, 62, 63, 69, 64, 65, 66, 67, 68, 70, 71, 72, 103, 104, 105] that learn a spatio-temporal model from training data, RED-PSM does not require ground-truth spatio-temporal training data, which is often not available or expensive to acquire. Thus RED-PSM offers the best of both worlds: a learned spatial model using readily available static training data, and unsupervised single-instance adaptation to the spatio-temporal measurement.
4. A novel and effective ADMM algorithm for the resulting new minimization problem enforces PSM as a hard constraint and incorporates the learned RED spatial regularizer.
5. The method is supported by theoretical analysis: we show convergence of the proposed PSM-based objective to a stationary point, and do so for a highly non-trivial learned regularizer. To date, the convergence of RED has been analyzed in the nonlinear measurement setting only for the phase retrieval problem [19]. Our analysis of RED-PSM is the first convergence result of RED in the bilinear (bi-convex), non-convex scenario. This convergence guarantee is of practical importance. First, it provides a mathematical justification to terminate the algorithm after sufficiently many iterations. Second, it eliminates concerns about the solution degrading after too many iterations as in the case of DIP-based methods.
6. Compared to a recent DIP-based [14] algorithm, RED-PSM achieves better reconstruction accuracy with orders of magnitude faster run times.
7. To improve and speed up the empirical convergence of RED-PSM, we use a particular fast projection-domain PSM initialization scheme [93, 94] for the spatial and temporal basis functions. The accelerated and reliable convergence with this initialization scheme is important for applications that require fast turn-around between acquisition and reconstruction.
8. A version of the approach with a patch-based regularizer is shown to provide almost equivalent reconstruction accuracy. This makes the proposed method conveniently scalable to high-resolution 3D or 4D settings.
An earlier and partial version of this work, missing, among other things, the detailed discussion of previous work, convergence analysis, and some of the experimental results, was presented at a conference [106].